OpenStudy (anonymous):

Consider an electron in the second excited state of a hydrogen atom. Which of the photon wavelengths can the atom absorb? a. 7.29 x 10^-8m b. 1.88x10^-6m The answer is B, I don't know how it was derived. I tried doing Rydberg equation but it just doesn't show.

OpenStudy (jfraser):

You'd need to know what level it is transitioning to, in order to use rydberg

OpenStudy (michele_laino):

Using Bohr's levels, and the conservation of energy, we get this expression: \[\Large \lambda = \frac{{8\pi }}{3}\frac{{\hbar c}}{{{E_0}}}\] where E0=13.6 eV nevertheless, I get this final result: \[\Large \lambda = 1.21 \cdot {10^{ - 7}}meters\]